erick's bank contains 66 quarters and dimes, of which Q are quarters. Find the total value of coins in dollars. He has $13.80
answer with steps?

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

Q+D=66 OR Q=66-D
25Q+10D=1380 NOW REPLACE (66-D) FOR Q & SOLVE FOR D.
25(66-D)+10D=1380
1650-25D+10D=1380
-15D=1380-1650
-15D=270
D=-270/-15
D=18 ANSWER FOR THE NUMBER OF DIMES.
Q=66-18=48 ANSWER FOR THE NUMBER OF QUARTERS.
PROOF
25*48+10*18=1380
1200+180=1380
1380=1380

Answer by edjones(8007)

q+d=66
.25q+.10d=13.80
multiply the bottom equation by 4 and subtract the top from the bottom.
q+.4d=55.2
q+d=66
-----------
-.6d=-10.8
divide -.6 into each side. d=18
replace d with 18 in the 1st equation. q+18=66
q=66-18
q=48
replace q & d with 48 & 18 respectively in the second equation.
.25*48+.1*18=13.8
12+1.8=13.8
Ans.= $12.00 in quarters. $1.80 in dimes.
Ed